OpenStudy (anonymous):

m(x)=X^3, f(x)=2x+4 First find F o m Your answer will be an algebraic expression, not a number. Next, find the first three iterates of f o m given xlower 0=-1

OpenStudy (anonymous):

F o M = f(m(x)) = 2x^3=4, just put in x^3 for x in your f function.

OpenStudy (anonymous):

does it matter if you have m o f first or do you get a different answer if you have the f first?

OpenStudy (anonymous):

m o f will get you a different answer (m(f(x)) = (2x+4)^3 = 8x^3 + 64). So, m o f =/= f o m. The order does matter when you do compositions of functions.